OpenStudy (anonymous):

Which of the following is a composite number? 51 17 29 89

OpenStudy (texaschic101):

51 is a composite number. It is not prime, because it has more then just 1 and itself as a factor. It also has 3 x 17.

OpenStudy (anonymous):

oooooo thank you!!!

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

how about this? Solve the equation (4x + 1)(3x - 2) = 0 {-4, 1.5} {-1/4, 2/3} {1/4, 2/3} {-4, 3}

OpenStudy (texaschic101):

(4x - 1)(3x - 2) = 0 4x - 1 = 0 4x = 1 x = 1/4 3x - 2 = 0 3x = 2 x = 2/3 answer (1/4,2/3)

OpenStudy (anonymous):

and what are the prime numbers in this list? Check all of the prime numbers in the following list. 3 11 27 19

OpenStudy (texaschic101):

take a look see... http://primes.utm.edu/lists/small/1000.txt

OpenStudy (anonymous):

that helped me thank you!!!!!

OpenStudy (anonymous):

What is the GCF of 40xy and 24x? Answer 8x 4xy 24x 40xy what is this texaschic???

OpenStudy (texaschic101):

8x

OpenStudy (texaschic101):

because 8 goes into 40 and 24 and x goes into xy and x

OpenStudy (anonymous):

Factor 18xy - 45yz using the GCF. 3y(6x - 15z) 3xy(6 - 15z) 9y(2x - 5z) 18yz(x - 2) is this the same?

OpenStudy (texaschic101):

18xy - 45yz) 9y(2x - 5z)

OpenStudy (texaschic101):

3y would factor but it is not the GCF

OpenStudy (anonymous):

thank you you have been such great help !! I appreciate it a lot!!

OpenStudy (texaschic101):

your welcome :)

OpenStudy (texaschic101):

do you understand how to do these ?

OpenStudy (texaschic101):

for instance, 6x + 2xy what goes into 6 and 2 ? what goes into x and xy ?

OpenStudy (anonymous):

6 gois in eith 2 I think and the rest I don't really now...

OpenStudy (texaschic101):

ok.....does 2 go into 6 and 2 ?

OpenStudy (texaschic101):

2 goes into 2 one time and 2 goes into 6 three times. x goes into xy so factored, it would be 2x(3 + 1y) or just 2x(3 + y)

OpenStudy (anonymous):

ooooo ok thank you anyways I have to go thanks for everything your the best see you lateeer :)

OpenStudy (texaschic101):

c ya later :)